Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this.-8x – 4y = -68x + 3y = 16
i have forgotten how this was done. can you help?

Answer by Alan3354(69443)

-8x – 4y = -68 --> 2x + y = 17
x + 3y = 16 --> x = 16 - 3y
--------------
Sub for x in the 1st eqn
2x + y = 17
2(16-3y) + y = 17
32 - 6y + y = 17
32 - 5y = 17
-5y = -15
y = 3
---------
Sub for y in either equation
x = 16 - 3y
x = 16 - 9
x = 7